OpenStudy (anonymous):

Find the area of a rectangular object which has a length of 5 square root of 8 inches and a width of 2 square root of 5 inches

OpenStudy (anonymous):

A = lw \[= 5\sqrt{8} \times 2\sqrt{5} = 10\sqrt{40} = 20\sqrt{10}\]